Sheet (Millions)

Income Statement (Millions)

Assets

Cash

1,290

Net Sales

51,600

Receivables

9,890

Operating Costs

48,246

Inventory

13,760

Depreciation

903

PP&E

18,060

EBIT

2,451

Total Assets

43,000

Interest

927

EBT

1,524

Liabilities and OE

Taxes

534

Accounts Payable

8,170

Net Income

990

Notes Payable

6,677

Wages Payable

4,073

Other Data

Long Term Bonds

8,815

Shares Outstanding (millions)

750

Total Equity

15,265

Common Dividends (millions)

$346.67

Total Liabilities and OE

43,000

Current Interest Rate

6.25%

Tax rate

35%

Stock Price

$35

Calculate P/E Ratio. Show all work
Answer #1

Solution:

P/E Ratio = Market price per share ÷ Earnings per share

Given, Market price per share = $35

Earnings per share = Earnings available to common shareholders ÷ shares outstanding

Earnings per share = $990millions ÷ 750millions

Earnings per share = 1.32

P/E Ratio = $35 ÷ 1.32 = 26.52

P/E Ratio = 26.52